OpenStudy (mayankdevnani):

If we toss one die 1000 times, how many times would we expect to get a 5? Answer about 56 about 100 about 168 about 200

OpenStudy (mayankdevnani):

@Hero @Luigi0210 @Miracrown

OpenStudy (anonymous):

is the answer 168

Directrix (directrix):

@mayankdevnani Is the die to be tossed a cubical, fair die?

OpenStudy (mayankdevnani):

fair die

OpenStudy (mayankdevnani):

i don't know the answer @No.name

OpenStudy (anonymous):

in one toss P(5)=1/6 1000 times toss=1/6*1000=?

OpenStudy (anonymous):

that's what , simple logic right @surjithayer

OpenStudy (anonymous):

so my answer is correct

OpenStudy (tkhunny):

@No.name Correct or not, simply posting the answer is not the right thing to do. Show your methods. Discuss your results. (See surjithayer, above) Ask important questions. (See Directrix, above).

OpenStudy (mayankdevnani):

thank you @surjithayer